OpenStudy (anonymous):

Find the equation of a line given the point and slope below. Arrange your answer in the form y = mx + b, where b is the constant. (5, 5) m = -3

OpenStudy (nurali):

m is the slope and (5,5) is a point on the line. Then for any other point (x,y) on the line you can write: (y-5)/(x-5) = -3 (definition of slope) then y-5=-3(x-5) = -3x +15. Solve for y to get it in the slope intercept form; y = -3x+10

OpenStudy (mathstudent55):

Here is another way of doing it. You want the form y = mx + b. You know m is -3, but you don't know b. That means you will have y = -3x + b You are given a point, (5, 5). Plug in the point for x and y and solve for b. 5 = -3(5) + b 5 = -15 + b 10 = b Now that you know b = 10, just replace it in the equation: y = -3x + 10

OpenStudy (anonymous):

can you help me with a few more?

OpenStudy (mathstudent55):

Sorry, it's very late for me. gtg

OpenStudy (nurali):

i will try.